OpenStudy (anonymous):

If 1 mole of XO3 contains the same number of atoms as 72 g of XH4, what is the relative atomic mass of X?

OpenStudy (mayankdevnani):

may i help you? @kryton1212

OpenStudy (anonymous):

but the answer is 86

OpenStudy (mayankdevnani):

If 1 mole of XO3 contains the same number of atoms as 72g of XH4, what is the relative atomic mass of X? 1 mole of XO3 contains 4 moles of atoms. 72g = 4 moles of atoms x 1molecule / 5 atoms = 0.8 moles of molecules = 0.8 moles XH4. 72g / 0.8moles = a molar mass of 90g/mole. Therefore the molar mass of X = 90-4 = 86g/mole for X.

OpenStudy (mayankdevnani):

got it!! @kryton1212

OpenStudy (anonymous):

got it, thanks

OpenStudy (mayankdevnani):

welcome